# An Embedding Theorem for tractor bundles, and an application in conformal pseudo-Riemannian geometry

### v1

@@ 4.2. Consequences of equivariance for stabilizer and isotropy
for $A\sqcup B=\{2,\ldots,p+1\}$, since the scalar product gives a nondegenerate pairing of each $\mathfrak{g}_{-\beta_{1}-\beta_{i}}$ with $\mathfrak{g}_{-\beta_{1}+\beta_{i}}$. The corresponding $\Sigma_{I}$ is $\{-\beta_{1}+\beta_{i}\}_{i\in A}$ union $\{-\beta_{1}-\beta_{i}\}_{i\in B}$. Now all values of $\rho$ are determined, and in particular $\rho(\beta_{i})\neq 0$ for all $i>1$.
Now $\iota(\mathfrak{h}_{-\alpha_{i^{\prime}}})\subseteq\mathfrak{g}_{-\beta_{1}}$ because $\rho(-\beta_{1}\pm\beta_{i})\neq-\alpha_{i^{\prime}}$ for any $i$. By (11),
$[\check{\mathfrak{h}}_{\alpha_{j^{\prime}}},\iota(\mathfrak{h}_{-\alpha_{i^{\prime}}-\alpha_{j^{\prime}}})]=[\iota(\mathfrak{h}_{\alpha_{j^{\prime}}}),\iota(\mathfrak{h}_{-\alpha_{i^{\prime}}-\alpha_{j^{\prime}}})]=\iota(\mathfrak{h}_{-\alpha_{i^{\prime}}})$
which means that for each $-\beta_{1}\pm\beta_{i}\in\Sigma_{I}$, there is some $j^{\prime}$ and there is $\beta^{\prime}\in\check{\Delta}(\alpha_{j^{\prime}})$ such that ±$\pm\beta_{i}+\beta^{\prime}=0$. Any such $\beta^{\prime}$ is a positive root, so $\Sigma_{I}=\{-\beta_{1}-\beta_{i}\ :\ i>1\}$, and $\iota(I)=\oplus_{j>1}\mathfrak{g}_{-\beta_{1}-\beta_{j}}$. Because $\mathfrak{h}_{-\alpha_{i^{\prime}}-\alpha_{j^{\prime}}}$ with $j^{\prime}\neq i^{\prime}$ is two-dimensional while each $\mathfrak{g}_{-\beta_{1}-\beta_{i}}$ is one-dimensional, there are at least two distinct $\beta_{j}$ with $\rho(\beta_{j})=\alpha_{j^{\prime}}$ for each $j^{\prime}\neq i^{\prime}$. There is moreover some $i\neq 1$ such that $\rho(\beta_{i})=\alpha_{i^{\prime}}$, and such that $\iota(\mathfrak{h}_{-2\alpha_{i^{\prime}}})=\mathfrak{g}_{-\beta_{1}-\beta_{i}}$. Since $p^{\prime}=(p+1)/2$, there are exactly two indices $j$ such that $\rho(\beta_{j})=\alpha_{j^{\prime}}$ for all $j^{\prime}$. The conclusion in this subcase follows.
Case 2: $\rho(2\beta_{1})=2\alpha_{i^{\prime}}+\alpha^{\prime}$ with $\alpha^{\prime}=\alpha_{j^{\prime}}$ or $\alpha_{j^{\prime}}+\alpha_{\ell^{\prime}}$ for some $j^{\prime}\neq i^{\prime}$, possibly with $j^{\prime}=\ell^{\prime}$. In this case $\overline{\mathfrak{h}}_{-\alpha_{i^{\prime}}}$ is isotropic, and belongs to an isotropic subspace together with $\oplus_{j^{\prime}\neq k^{\prime}}\overline{\mathfrak{h}}_{-\alpha_{i^{\prime}}-\alpha_{k^{\prime}}}$; moreover $\overline{\mathfrak{h}}_{-\alpha_{i^{\prime}}}$ has dimension $2$. As in the previous subcase, any $\overline{\mathfrak{h}}_{\alpha_{i^{\prime}}-\alpha_{k^{\prime}}}$ for $k^{\prime}\neq i^{\prime}$ could be added to make a larger isotropic subspace, so these root spaces are contained in $\mathfrak{h}^{0}$, and therefore also $\mathfrak{h}_{\alpha_{i^{\prime}}}\subset\mathfrak{h}^{0}$.
@@
As before, $\iota(J)$ is contained in $\mathfrak{g}_{-}$ and $\iota(I)$ is a maximal isotropic subspace. The subset $\Sigma_{I}\subset\Sigma$ is defined as above, the roots in $\Sigma$ for which $\iota(I)$ has nontrival projection on the corresponding root space. Since $\rho(-\beta_{1})$ is not a root occuring in $I$, then $-\beta_{1}\notin\Sigma_{I}$. Thus $\iota(I)$ has the same form as (12). By dimension considerations, the restriction of $\rho$ to $\Sigma_{I}$ is a 2-to-1 map to the roots of $\Delta_{H}$ occuring in $I$, except to $-2\alpha_{i^{\prime}}$, onto which it is 1-to-1. The values of $\rho(\beta_{i})$ are thus determined for all $i$.
@@ 4.8.2. Isotropic tractor solutions
The following consequences of balanced linear isotropy are already known in the conformal Lorentzian case (see [13, Lem 6.5, Thm 1.3(2)], [23, Sec 4]), and generalize to higher signature as follows.
@@
If $(M,g)$ is analytic and simply connected, then the solutions in $\mathcal{I}_{U}$ extend to a subspace $\mathcal{I}\subset\mathcal{S}$, and $X$ extends to an isotropic conformal vector field on $M$.
@@
In this frame, the Weyl tensor $W_{x_{0}}$ is a homomorphism $\hat{W}_{\hat{x}_{0}}:\wedge^{2}{\bf R}^{p,q}\rightarrow\mathfrak{so}(p,q)$, which is $\{h^{t}\}$-equivariant. The eigenvalues of Ad $\mbox{Ad }h^{t}$ on $\mathfrak{so}(p,q)$ are $e^{t},1,$ and $e^{-t},$ and the eigenspace for $e^{-t}$ is precisely the nilpotent subalgebra ${\mathfrak{n}}(e_{n})$. It follows that $\hat{W}_{\hat{x}_{0}}(u,v)\in{\mathfrak{n}}(e_{n})$, for all $u,v\in{\bf R}^{p,q}$.
The flow $\{h^{t}\}$ is stable for $t\rightarrow\infty$ in the sense of [13, Def 4.2]. It follows that $\{h^{t}\}$ remains a holonomy flow for points nearby $x_{0}$, meaning that there is a neighborhood $U$ of $x_{0}$ with a section $\hat{U}\subset\hat{M}$ containing $\hat{x}_{0}$ such that $\varphi^{t}_{Y}.\hat{x}.h^{-t}$ is bounded in $\hat{M}$ as $t\rightarrow\infty$ for all $\hat{x}\in\hat{U}$ (see [7, Prop 2.8]); moreover, the section $\hat{U}$ can be taken to be $C^{\omega}$. For any $\hat{x}\in\hat{U}$, the Weyl curvature $\hat{W}_{\hat{x}}$ must remain bounded—though not necessarily invariant—under the action of $\{h^{t}\}$. It follows that $\hat{W}_{\hat{x}}(u,v)\in{\mathfrak{n}}(e_{n})$ for all $u,v\in{\bf R}^{p,q}$, $\hat{x}\in\hat{U}$.
Because it is $\{\varphi^{t}_{Y}\}$-invariant, the full Cartan curvature $\kappa_{\hat{x}}$ is bounded under the action of the holonomy $\{h^{t}\}$. The ${\mathfrak{p}}_{+}$-component of $\kappa_{\hat{x}}$ is in $\wedge^{2}{\bf R}^{p,q*}\otimes{\bf R}^{p,q*}$, which has no stable elements under $\{h^{t}\}$. Consequently, $\kappa_{\hat{x}}=\hat{W}_{\hat{x}}$ for all $\hat{x}\in\hat{U}$.
Let $X\in\mathcal{X}(U)$ be the isotropic vector field with values $X_{x}$ equal $e_{n}$ in the frame corresponding to $\hat{x}\in\hat{U}$, for each $x\in U$. It determines a reduction $\mathcal{R}_{U}$ of $\hat{M}_{U}$ to a subgroup $Q\cong Q_{0}\ltimes Q_{1}<P$, where $Q_{0}<\mbox{CO}(p,q)$ is the stabilizer of ${\bf R}e_{n}$ and $Q_{1}<P^{+}\cong{\bf R}^{p,q*}$ is the annihilator of $e_{n}^{\perp}$ (compare [23, Sec 4]). Because $Q$ normalizes ${\mathfrak{n}}(e_{n})$, the Cartan curvature retains the properties of being equal to the Weyl curvature and of having values in ${\mathfrak{n}}(e_{n})$ on all of $\mathcal{R}_{U}$. It follows from the formula, for $U,V\in\mathcal{X}(\hat{M})$,
@@
that the distribution $\mathcal{D}=\omega^{-1}({\bf R}^{p,q}+{\mathfrak{n}}(e_{n}))$ is integrable in restriction to $\mathcal{R}_{U}$ (see for example [23, Sec 4]).
Let $\mathcal{L}_{0}$ be the leaf through $\hat{x}_{0}$ tangent to $\mathcal{D}$. It is a reduction to $U(e_{n})$, the unipotent subgroup of $G_{0}$ with Lie algebra ${\mathfrak{n}}(e_{n})$. This reduction corresponds to a metric $g_{0}\in[\left.g\right|_{U}]$ which is Ricci-flat and for which $X$ is parallel. We have proved (1)-(3).
@@
When $(M,[g])$ is real-analytic, then any local solution lifts and extends to a global solution on the pullback of $\mathcal{V}$ to the universal cover of ${M}$, and the same holds for local conformal vector fields [2]. Under the assumption that $M$ is simply connected, any local solutions globalize, as does the local conformal vector field $X$. Then the solutions $\mathcal{I}_{U}$ extend to global solutions $\mathcal{I}\subset\mathcal{S}$, and the extension of $X$ is everywhere isotropic by analyticity, as it is isotropic on $U$. ∎
@@
Now $\hat{x}_{0}$ belongs to the reduction of $\hat{M}$ to $U(e_{n})$ constructed in the proof of Proposition 4.13 above, corresponding to the Ricci-flat metric $g_{0}$. The trivialization $\iota_{\hat{x}_{0}}:\mathcal{V}_{x_{0}}\rightarrow{\bf V}$ therefore identifies the subspaces $(\mathcal{V}_{i})_{x_{0}}$ of the decomposition determined by $g_{0}$ with ${\bf V}_{i}$ for $i=1,0,-1$. The constant sections of $\mathcal{V}_{-1}$, belonging to $\mathcal{I}$, corresponding to metrics homothetic to $g_{0}$, therefore evaluate under $\iota_{\hat{x}_{0}}$ to elements of ${\bf V}_{-1}={\bf R}E_{n+1}$.

### v2

@@ 4.2. Consequences of equivariance for stabilizer and isotropy
for $A\sqcup B=\{2,\ldots,p+1\}$ The corresponding $\Sigma_{I}$ is $\{-\beta_{1}+\beta_{i}\}_{i\in A}$ union $\{-\beta_{1}-\beta_{i}\}_{i\in B}$. Thus all values of $\rho$ are determined; in particular $\rho(\beta_{i})\neq 0$ for all $i>1$.
Now $\overline{\iota(\mathfrak{h}_{-\alpha_{i^{\prime}}})}\subseteq\overline{\mathfrak{g}}_{-\beta_{1}}$ because $\rho(-\beta_{1}\pm\beta_{i})\neq-\alpha_{i^{\prime}}$ for any $i$; in particular $p<q$ in this case. By (11),
$\check{\mathfrak{h}}_{\alpha_{j^{\prime}}},\iota(\mathfrak{h}_{-\alpha_{i^{\prime}}-\alpha_{j^{\prime}}})]=[\iota(\mathfrak{h}_{\alpha_{j^{\prime}}}),\iota(\mathfrak{h}_{-\alpha_{i^{\prime}}-\alpha_{j^{\prime}}})]=\iota(\mathfrak{h}_{-\alpha_{i^{\prime}}})$
which means that for each $\beta_{1}\pm\beta_{i}\in\Sigma_{I}$, there is some $j^{\prime}$ and there is $\beta^{\prime}\in\check{\Delta}(\alpha_{j^{\prime}})$ such that $\pm\beta_{i}+\beta^{\prime}=0$. As $\beta^{\prime}$ is a positive root, we conclude $\Sigma_{I}=\{-\beta_{1}-\beta_{i}\ :\ i>1\}$, and $\overline{\iota(I)}=\oplus_{j>1}\overline{\mathfrak{g}}_{-\beta_{1}-\beta_{j}}$. Because $\overline{\mathfrak{h}}_{-\alpha_{i^{\prime}}-\alpha_{j^{\prime}}}$ with $j^{\prime}\neq i^{\prime}$ is two-dimensional while each $\overline{\mathfrak{g}}_{-\beta_{1}-\beta_{i}}$ is one-dimensional, there are at least two distinct $\beta_{j}$ with $\rho(\beta_{j})=\alpha_{j^{\prime}}$ for each $j^{\prime}\neq i^{\prime}$. There is moreover some $i\neq 1$ such that $\rho(\beta_{i})=\alpha_{i^{\prime}}$, and such that $\overline{\iota(\mathfrak{h}_{-2\alpha_{i^{\prime}}})}=\overline{\mathfrak{g}}_{-\beta_{1}-\beta_{i}}$. Since $p^{\prime}=(p+1)/2$, there are exactly two indices $j$ such that $\rho(\beta_{j})=\alpha_{j^{\prime}}$ for all $j^{\prime}$. The conclusion in this subcase follows.
Case 2: $\rho(2\beta_{1})=2\alpha_{i^{\prime}}+\alpha^{\prime}$ with $\alpha^{\prime}=\alpha_{j^{\prime}}$ or $\alpha_{j^{\prime}}+\alpha_{\ell^{\prime}}$ for some $j^{\prime}\neq i^{\prime}$, possibly with $j^{\prime}=\ell^{\prime}$. In this case $\overline{\mathfrak{h}}_{-\alpha_{i^{\prime}}}$ is $[s]$-isotropic, and belongs to an isotropic subspace together with $\oplus_{j^{\prime}\neq k^{\prime}}\overline{\mathfrak{h}}_{-\alpha_{i^{\prime}}-\alpha_{k^{\prime}}}$. By Lemma 4.5 (2), each $\mathfrak{h}_{-\alpha_{i^{\prime}}-\alpha_{k^{\prime}}}$ has trivial intersection with $\mathfrak{h}^{0}$; also $\overline{\mathfrak{h}}_{-\alpha_{i^{\prime}}}$ has dimension $2$. As in the previous subcase, any $\overline{\mathfrak{h}}_{\alpha_{i^{\prime}}-\alpha_{k^{\prime}}}$ for $k^{\prime}\neq i^{\prime}$ could be added to make a larger isotropic subspace, so these root spaces are contained in $\mathfrak{h}^{0}$, and therefore also $\mathfrak{h}_{\alpha_{i^{\prime}}}\subset\mathfrak{h}^{0}$.
@@
As before, $\overline{\iota(I)}$ is a maximal isotropic subspace of $\mathfrak{g}/{\mathfrak{p}}\cong{\bf R}^{p,q}$. The subset $\Sigma_{I}\subset\Sigma$ is defined as above. Since $\rho(-\beta_{1})$ is not a root occuring in $I$, then $\beta_{1}\notin\Sigma_{I}$. Thus $\overline{\iota(I)}$ has the same form as (13). By dimension considerations, the restriction of $\rho$ to $\Sigma_{I}$ is a 2-to-1 map to the roots of $\Delta_{H}$ occuring in $I$, except to $2\alpha_{i^{\prime}}$, onto which it is 1-to-1. The values of $\rho(\beta_{i})$ are thus determined for all $i$.
@@ 4.8.2. Isotropic tractor solutions
The following consequences of balanced linear isotropy are already known in the conformal Lorentzian case (see [14, Lem 6.5, Thm 1.3(2)], [23, Sec 4]), and generalize to higher signature as follows.
@@
If $(M,g)$ is analytic and simply connected, then the solutions in $\mathcal{I}_{U}$ extend to $M$, comprising a subspace $\mathcal{I}\subset\mathcal{S}$, and $X$ extends to an isotropic conformal vector field on $M$.
@@
In this frame, the Weyl tensor $W_{x_{0}}$ is a homomorphism $\hat{W}_{\hat{x}_{0}}:\wedge^{2}{\bf R}^{p,q}\rightarrow\mathfrak{so}(p,q)$, which is $\{h^{t}\}$-equivariant. The eigenvalues of $\mbox{Ad }h^{t}$ on $\mathfrak{so}(p,q)$ are $e^{t},1,$ and $e^{-t},$ and the eigenspace for $e^{-t}$ is precisely the nilpotent subalgebra ${\mathfrak{n}}(e_{n})$. It follows that $\hat{W}_{\hat{x}_{0}}(u,v)\in{\mathfrak{n}}(e_{n})$, for all $u,v\in{\bf R}^{p,q}$.
The flow $\{h^{t}\}$ is stable for $t\rightarrow\infty$ in the sense of [14, Def 4.2]. It follows that $\{h^{t}\}$ remains a holonomy flow for points nearby $x_{0}$, meaning that there is a neighborhood $U$ of $x_{0}$ with a section $\hat{U}\subset\hat{M}$ containing $\hat{x}_{0}$ such that $\varphi^{t}_{Y}.\hat{x}.h^{-t}$ is bounded in $\hat{M}$ as $t\rightarrow\infty$ for all $\hat{x}\in\hat{U}$ (see [8, Prop 2.8]); moreover, the section $\hat{U}$ can be taken to be $C^{\omega}$. For any $\hat{x}\in\hat{U}$, the Weyl curvature $\hat{W}_{\hat{x}}$ must remain bounded—though not necessarily invariant—under the action of $\{h^{t}\}$. It follows that $\hat{W}_{\hat{x}}(u,v)\in{\mathfrak{n}}(e_{n})$ for all $u,v\in{\bf R}^{p,q}$, $\hat{x}\in\hat{U}$.
[…]
Let $X\in\mathcal{X}(U)$ be the isotropic vector field with values $X_{x}$ equal $e_{n}$ in the frame corresponding to $\hat{x}\in\hat{U}$, for each $x\in U$. It determines a reduction $\mathcal{R}_{U}$ of $\hat{M}_{U}$ to a subgroup $Q\cong Q_{0}\ltimes Q_{1}<P$, where $Q_{0}<\mbox{CO}(p,q)$ is the stabilizer of ${\bf R}e_{n}$ and $Q_{1}<P^{+}\cong{\bf R}^{p,q*}$ is the annihilator of $e_{n}^{\perp}$, as in the conformal Lorentzian case [23, Sec 4]. First the reduction $\mathcal{R}^{\prime}_{U}\subset\hat{M}^{\prime}=\hat{M}/P^{+}$ is defined to be the conformal frames which identify $X$ with ${\bf R}e_{n}$; it is an analytic reduction to $Q_{0}$. Then $\mathcal{R}_{U}$ is defined as the set of $\hat{x}$ such that the projection $\pi^{\prime}(\hat{x})=\hat{x}^{\prime}\in\mathcal{R}^{\prime}_{U}$ and $\pi^{\prime}_{*}(\omega_{\hat{x}}^{-1}(\mathfrak{g}_{-1}))\subset T_{\hat{x}^{\prime}}\mathcal{R}^{\prime}_{U}$. The key observation is that
${\mathfrak{q}}_{1}=\{\xi\in{\mathfrak{p}}^{+}\ :\ [\xi,\mathfrak{g}_{-1}]\subset{\mathfrak{q}}_{0}\},$
from which
$Q_{1}=\{g\in P^{+}\ :\ (\mbox{Ad }g)(\mathfrak{g}_{-1}+{\mathfrak{q}}_{0})\equiv(\mathfrak{g}_{-1}+{\mathfrak{q}}_{0})\ \mbox{mod }{\mathfrak{p}}^{+}\}.$
Now it follows that $\mathcal{R}_{U}$ is a reduction of $\hat{M}$ to $Q$; it is also analytic.
[…]
@@
that the distribution $\mathcal{D}=\omega^{-1}({\bf R}^{p,q}+{\mathfrak{n}}(e_{n}))$ is integrable in restriction to $\mathcal{R}_{U}$ (see for example [23, Sec 4]). Let $\mathcal{L}_{0}$ be the leaf through $\hat{x}_{0}$ tangent to $\mathcal{D}$. It is a reduction to $U(e_{n})$, the unipotent subgroup of $G_{0}$ with Lie algebra ${\mathfrak{n}}(e_{n})$. This reduction corresponds to a metric $g_{0}\in[\left.g\right|_{U}]$ which is Ricci-flat and for which $X$ is parallel. We have proved (1)-(3).
@@
When $(M,[g])$ is real-analytic, then any local solution lifts and extends to a global solution on the pullback of $\mathcal{V}$ to the universal cover of M{M}, and the same holds for local conformal vector fields [2]. Under the assumption that $M$ is simply connected, any local solutions globalize, as does the local conformal vector field $X$. Then the solutions $\mathcal{I}_{U}$ extend to global solutions $\mathcal{I}\subset\mathcal{S}$, and the extension of $X$ is everywhere isotropic by analyticity, as it is isotropic on $U$. ∎
@@
Now $\hat{x}_{0}$ belongs to the reduction of $\hat{M}$ to $U(e_{n})$ constructed in the proof of Proposition 4.13 above, corresponding to the Ricci-flat metric $g_{0}$. The trivialization $\iota_{\hat{x}_{0}}:\mathcal{V}_{x_{0}}\rightarrow{\bf V}$ therefore identifies the subspaces $(\mathcal{V}_{i})_{x_{0}}$ of the decomposition determined by $g_{0}$ with ${\bf V}_{i}$ for i=1,0,1i=1,0,-1. The constant sections of $\mathcal{V}_{-1}$, belonging to $\mathcal{I}$, corresponding to metrics homothetic to $g_{0}$, therefore evaluate under $\iota_{\hat{x}_{0}}$ to elements of ${\bf V}_{-1}={\bf R}E_{n+1}$.